# ReGNet: Reciprocal Space-Aware Long-Range Modeling for Crystalline Property Prediction

### v2

@@ 5.2 Ablation Studies
Table 4: Ablation studies of reciprocal space and model depth in test MAE on the JARVIS dataset.
Method Reciprocal Blocks Form. Energy Bandgap(OPT) $E_{total}$ Bandgap(MBJ) $E_{hull}$ ReGNet w/o Reciprocal 3 29.6 0.135 32.5 0.295 68.8 ReGNet 3 27.8 0.127 27.2 0.242 52.9 ReGNet 4 27.0 0.126 26.8 0.239 48.0
In this section, we examine the importance of incorporating reciprocal space and analyze the effect of the number of ReGNet blocks through ablation studies on the JARVIS dataset. Specifically, we evaluate a variant of ReGNet with the reciprocal space component removed. As shown in Table 4, including reciprocal space information leads to improved performance. Furthermore, increasing the number of ReGNet blocks from 3 to 4 yields additional gains on the JARVIS benchmark. Additional ablation results are provided in Section B.5.
@@ 5.3 Multi-property prediction
ReGNet-MT achieves promising performance with much higher efficiency. On the Materials Project dataset, we apply two ReGNet-MT models separately for predicting two energy properties and two mechanical properties. It achieves comparable performance as shown in Table 13 in the Appendix. For the JARVIS dataset, ReGNet-MT on 5-property prediction delivers the best results for band gap (OPT) and band gap (MBJ), outperforming all single-property prediction models, while maintaining comparable performance on the remaining tasks to Matformer as shown in Table 12, which benefits from positive transfer as discussed in Section C.5. Besides, its efficiency is discussed in Section 5.4.
@@ 5.4 Model Efficiency Comparison
This section highlights the efficiency of ReGNet in handling both single-task and multi-task scenarios. As shown in Table 5, we evaluate the computational efficiency of ReGNet and ReGNet-MT by comparing their model parameters and training time per epoch against ALIGNN, Matformer, PotNet, and ComFormer on the JARVIS formation energy task. We use 3 and 4 to indicate the number of blocks in ReGNet. ReGNet demonstrates superior computational efficiency, with ReGNet(3) achieving the fastest training time per epoch among all models. Notably, in the multi-task setting, ReGNet-MT predicts five properties simultaneously. The training time does not even double that of single-task ReGNet, and the model’s parameter size is smaller than that of single-task eComFormer. It is also important to note that PotNet reduces its training time by precomputing infinite potential summations for the training data, which is not employed in ReGNet.
Table 5: Comparison of model efficiency on JARVIS formation energy prediction. We include training time per epoch, 1-task model parameter size, and 5-task model parameter size. N/A denotes that ReGNet-MT is not designed for single-property prediction.
Method Type Time/Epoch 1-task Model Para. 5-task Model Para.
ALIGNN GNN 261 s 4.0 M 20.0 M
Matformer Transformer 51 s 2.9 M 14.5 M
PotNet GNN 34 s 1.8 M 9.0 M
iComFormer Transformer 62 s 5.0 M 25.0 M
eComFormer Transformer 92 s 12.4 M 62.0 M
ReGNet(3) GNN 31 s 3.3 M 16.5 M
ReGNet(4) GNN 38 s 4.3 M 21.5 M
ReGNet-MT GNN 60 s N/A 9.5 M
6 Limitation
One limitation of our study lies in using geometric GNNs for short-range modeling to capture local features. While effective, recent advancements in Transformer-based models (Yan et al., 2022, 2024) present opportunities for further improvement by incorporating reciprocal space information to enhance the structural representation of materials. Besides, multi-property prediction remains an area worthy of further exploration, as current extensions only provide an initial attempt. Despite these limitations, our study provides a new perspective on modeling crystal materials by leveraging long-range information from reciprocal space alongside graph-based short-range information.
7 Conclusion
In this paper, we introduce a novel model, ReGNet, designed to capture the reciprocal space-based long-range information and geometric short-range information in crystal property prediction. Besides, its multi-task variant with the help of MoE has achieved promising results in multi-property prediction. Empirically, ReGNet achieves significant performance improvement on three widely used crystal benchmarks. We hope this efficient model provides a valuable perspective to both machine learning and materials science, promoting further interdisciplinary research and encouraging its application across diverse material systems as a foundational component.
@@ A.1 Relations Between Real and Reciprocal Lattice Vectors
In materials science, crystals exhibit translational symmetry, where atoms are periodically arranged in 3D space, forming a direct lattice in real space. Each lattice point corresponds to a repeating unit in the crystal structure, described by primitive lattice vectors $\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}$. Reciprocal space, on the other hand, represents the spatial frequencies associated with this periodic arrangement. It forms a dual lattice, mathematically related to the real lattice through reciprocal lattice vectors $\bm{b}_{1},\bm{b}_{2},\bm{b}_{3}$. $V$ is the volume of the parallelepiped spanned by the three primitive translation vectors of the original Bravais lattice. It remains invariant under periodic permutations of the indices (Economou, 2010).
@@ A.2 Diffraction Methods
Electron diffraction is widely used in transmission electron microscopy (TEM) to analyze materials. For instance, single crystals produce discrete spot patterns, polycrystals form concentric rings, and amorphous materials show diffuse rings (Egerton et al., 2005). These patterns enable crystallinity and orientation analysis, with applications in semiconductors, metals, and oxides. Figure 3 illustrates a typical diffraction pattern. By satisfying the Bragg condition, elastically scattered electrons produce high-intensity spots, with the angular relationship between transmitted and diffracted beams revealing the crystal structure (Poeppelmeier, 2023).
@@ B.1 Dataset
JARVIS-DFT Formation Energy 55722
Bandgap (OPT) 55722
Total Energy 55722
Bandgap (MBJ) 18171
$E_{\text{hull}}$ 55370
Materials Project Formation Energy 69239
Band Gap 69239
Bulk Moduli 5450
Shear Moduli 5450
MatBench e_form 132752
@@
JARVIS Dataset
The JARVIS dataset, proposed by Choudhary et al. (2020), contains 55723 materials and serves as a benchmark for crystal property prediction. Following the experimental protocols of Yan et al. (2024), we evaluate our methods on five regression tasks: formation energy, total energy, bandgap (using both OPT and MBJ methods), and energy above the convex hull (Ehull). The training, validation, and testing splits for formation energy, total energy, and OPT bandgap consist of 44578, 5572, and 5572 crystals, respectively. The splits for Ehull include 44296, 5537, and 5537 crystals, while the MBJ bandgap tasks consist of 14537, 1817, and 1817 crystals. The MBJ functional is considered more accurate for bandgap calculations compared to OPT, and both are utilized in this study. Notably, 18865 of the dataset’s crystal structures have been experimentally observed, adding robustness to its use in predictive modeling.
The Materials Project Dataset
The Materials Project dataset, introduced by Chen et al. (2019), is a comprehensive collection of 69239 materials, widely utilized in crystal property prediction studies. We adopt the data splits in Yan et al. (2024) to ensure consistency and fair comparisons with prior methods. Specifically, for formation energy and bandgap prediction tasks, the dataset comprises 60000, 5000, and 4239 crystals for training, validation, and testing, respectively. For bulk modulus and shear modulus predictions, the splits include 4664, 393, and 393 crystals. Notably, one validation sample in the shear modulus task is excluded due to a negative GPa value, reflecting an underlying unstable or metastable crystal structure. Among the included crystals, 30084 have been experimentally observed, further highlighting the dataset’s reliability for studying material properties.
MatBench Dataset
The Matbench dataset Dunn et al. (2020) is a benchmark suite designed for evaluating machine learning models on materials science tasks, offering standardized splits and diverse property predictions. In this work, we focus on two representative regression tasks from Matbench: the large-scale formation energy prediction task (e_form) comprising 132752 crystal structures, and the smaller-scale 2D exfoliation energy task (jdft2d) with 636 samples. These tasks assess the model’s ability to generalize across both abundant and limited data regimes.
@@ B.2 Settings of Model Embeddings
Node feature is embedded into a CGCNN (Xie & Grossman, 2018) embedding vector of length 92 based on atomic number, and then mapped to a 256-dimensional initial short-range node feature ${h}^{0}_{\text{local}}$ by a linear transformation. The initial global node feature ${h}^{0}_{\text{global}}$ is obtained by a linear transformation on ${h}^{0}_{\text{local}}$ and follows an activation function. The invariant edge feature $||\mathbf{p}_{i}-\mathbf{p}_{j}||_{2}$ is inversely scaled to $-0.75/||\mathbf{p}_{i}-\mathbf{p}_{j}||_{2}$ based on Lin et al. (2023), and then expanded into a 256-dimensional vector by RBF kernels with 256 center values from -4.0 to 4.0; after that, we transfer the 256-dimensional vector to initial edge feature ${v}_{ij}^{e}$ through a linear transformation followed by a SoftPlus activation.
@@ B.3 Geometric GNN Information
[…]
@@
B.4 Hyperparameter Settings of ReGNet on Different Tasks
In this subsection, we share the detailed hyperparameter settings of ReGNet and ReGNet-MT for different tasks in JARVIS, the Materials Project, and MatBench crystal datasets. We use a different number of blocks for different tasks, and higher performance is expected if hyperparameters are tuned specifically for each task.
Table 7: Model settings of ReGNet for the JARVIS dataset.
@@ B.4 Hyperparameter Settings of ReGNet on Different Tasks
Table 8: Model settings of ReGNet for the Materials Project dataset.
@@
Table 9: Model settings of ReGNet-MT for JARVIS and the Materials Project datasets.
@@
Table 10: Model settings of ReGNet for MatBench.
@@
JARVIS. We show the model settings of ReGNet on the JARVIS dataset in Table 7. We train both models using the Adam (Kingma, 2014) optimizer with weight decay of 1e–5, Onecycle scheduler (Smith & Topin, 2019), and for a duration of 500 training epochs. The batch size is standardized at 64, and the models are trained using L1 loss function. The effectiveness of the model is quantitatively measured using the mean absolute error (MAE). The number of neighbors indicates the $k$-th nearest distance we use as the radius for node $i$.
The Materials Project. We show the model settings of ReGNet on the Materials Project dataset in Table 8. We train both models using the Adam (Kingma, 2014) optimizer with weight decay of 1e–5, and the Onecycle scheduler (Smith & Topin, 2019). The batch size is standardized at 64, and the models are trained using L1 loss function. The effectiveness of the model is quantitatively measured using the mean absolute error (MAE). The number of training epochs for shear moduli on ReGNet is 300, while for the other three properties it is 500.
MatBench. We show the model settings of ReGNet on the MatBench dataset in Table 10. All models are trained for 500 epochs using the Adam (Kingma, 2014) optimizer with MAE loss. The effectiveness of the models is quantitatively measured using the mean absolute error (MAE) and root mean square error (RMSE).
Figure 4: Efficiency comparison of training time per epoch, with each method’s training time scaled relative to ReGNet(3). 3 and 4 indicate the number of ReGNet blocks.
B.5 Detailed Ablation Studies
Table 11: Ablation study of ReGNet with varying model depth, reported in terms of test MAE on the JARVIS dataset. The numbers 3, 4, and 5 denote the number of ReGNet blocks used. The best results are shown in bold.
@@ B.5 Detailed Ablation Studies
ReGNet(3) 27.8 0.127 27.2 0.242 52.9
ReGNet(4) 27.0 0.126 26.8 0.239 48.0
ReGNet(5) 27.7 0.127 27.2 0.240 43.0
@@
We further investigate the effect of model depth by varying the number of ReGNet blocks from 3 to 5. As shown in Table 11, increasing the number of blocks from 3 to 4 consistently improves performance across all tasks. Performance on $E_{\text{hull}}$ continues to improve with five blocks, suggesting that deeper message passing may help capture the subtle energy differences relevant to thermodynamic stability. However, for other properties on the JARVIS dataset, adding a fifth block offers no additional benefit and leads to slight performance degradation. This suggests that while increased depth enhances representational capacity, it may also introduce overfitting or optimization difficulties, especially when training data is limited. To balance predictive accuracy and computational cost, we adopt three or four ReGNet blocks as the default configuration.
B.6 Training Time Scale Efficiency Analysis
Beyond the superior modeling capacity for crystalline materials, our ReGNet is faster and more efficient than previous works. To demonstrate the efficiency of ReGNet, we compare ReGNet and its multi-task variant ReGNet-MT with iComFormer, eComFormer, PotNet, Matformer, and ALIGNN. Evaluation is in terms of training time per epoch on the task of JARVIS formation energy prediction. From Figure 4, ReGNet demonstrates superior computational efficiency. Specifically, ReGNet(3) outperforms iComFormer by 1.77$\times$, eComFormer by 2.96$\times$, PotNet by 1.08$\times$, Matformer by 1.65$\times$, and ALIGNN by 8.41$\times$ speedup. Finally, the time for ReGNet-MT predicting 5 properties does not even double. These results highlight the effectiveness of ReGNet in reducing computational overhead while maintaining performance in prediction, as previously mentioned.
Appendix C Multi-task Learning
C.1 Motivation
While our novel usage of reciprocal space representations effectively captures long-range interactions and global periodicity, it still adheres to the traditional approach of predicting properties individually Schütt et al. (2017); Choudhary & DeCost (2021); Yan et al. (2024). This approach becomes computationally and memory inefficient when repeatedly applied to multi-property predictions. To address this limitation, Multi-Task Learning (MTL) offers a promising solution by enabling the prediction of multiple properties within a single model Evgeniou & Pontil (2004). Additionally, MTL can leverage positive transfer, where knowledge and representations learned from one task can benefit other tasks. This advantage is particularly relevant for materials, as similar properties (e.g., OPT bandgap and MBJ bandgap) are often highly correlated and share similar structural features. However, to our knowledge, few studies have specifically investigated multi-property prediction for materials.
C.2 Related Works on Multi-property prediction

### v3

@@ 5.2 Ablation Studies
We further examine the importance of incorporating reciprocal space and analyze the effect of the number of ReciNet blocks through ablation studies on the JARVIS dataset. We evaluate a variant of ReciNet with the reciprocal space component removed. As shown in Table˜4, including reciprocal space information leads to improved performance. Furthermore, adding one more reciprocal block leads to improved performance. Additional ablation results are provided in Section˜B.5.
@@ 5.3 Multi-property prediction
We include ReciNet-MT as an extension to demonstrate the scalability of our design in multi-property settings, with detailed descriptions and experiments in Appendix˜C. On the Materials Project dataset, we apply two ReciNet-MT models separately for predicting two energy properties and two mechanical properties. It achieves comparable performance as shown in Table˜14 in the Appendix. For the JARVIS dataset, ReciNet-MT on 5-property prediction delivers the best results for band gap (OPT) and band gap (MBJ) while maintaining comparable performance on the remaining tasks to Matformer as shown in Table˜13. The observed improvement can be attributed to positive transfer as discussed in Section˜C.5. Besides, its efficiency is discussed in Section˜5.4.
@@ 5.4 Model Efficiency Comparison
Table˜5 reports the efficiency of top-performing architectures in terms of runtime and model size. ReciNet attains superior accuracy while requiring only 12% of the training time of CrystalFramer, 42% of iComFormer, and 66% of Crystalformer. Although Crystalformer and CrystalFramer incur low per-epoch costs, they require a large number of training epochs (800 and 2,000, respectively) to converge. By contrast, ReciNet combines competitive per-epoch and inference efficiency with a significantly reduced epoch budget. In the multi-task setting, ReciNet-MT predicts five properties simultaneously with less than twice the training time of single-task ReciNet, while also using fewer parameters than single-task eComFormer. It is also important to note that PotNet reduces its training time by precomputing potential summations, which is shown in inference time as domination.
Table 5: Efficiency comparison. Comparison of model efficiency on JARVIS formation energy prediction. We include training time per epoch, training epochs, total training time, and average inference time per material. We also show the number of parameters for 1 and 5 tasks. N/A denotes that ReciNet-MT is not designed for single-property prediction.
Model Type Time/Epoch Epochs Total Inference 1-task Para. 5-task Para.
PotNet (Lin et al., 2023) GNN 43 s 500 6.0 h 313 ms 1.8 M 9.0 M
Matformer (Yan et al., 2022) Transformer 60 s 500 8.3 h 20.4 ms 2.9 M 14.5 M
iComFormer (Yan et al., 2024) Transformer 59 s 700 11.5 h 54.8 ms 5.0 M 25.0 M
Crystalformer (Taniai et al., 2024) Transformer 32 s 800 7.2 h 6.6 ms 853 K 4.27 M
CrystalFramer (Ito et al., 2025) Transformer 74 s 2000 41.2 h 16.8 ms 952 K 4.76 M
ReciNet GNN 35 s 500 4.8 h 8.2 ms 3.3 M 16.5 M
ReciNet-MT GNN 69 s 500 9.6 h 16.2ms N/A 9.5 M
6 Conclusion, Limitations, and Future Work
In this work, we introduce ReciNet, a novel model designed to capture both reciprocal space-based long-range and geometric short-range information for crystal property prediction. Empirically, ReciNet achieves substantial performance improvements on three widely used crystal benchmarks, while its multi-task variant with mixture-of-experts (MoE) demonstrates promising results in multi-property prediction. There are several limitations for future enhancement. Firstly, our short-range modeling with geometric GNNs, while effective, could be improved by incorporating recent advancements in Transformer-based models to enhance material representations. Secondly, our work in multi-property prediction represents an initial step that warrants further exploration. Overall, we encourage the adoption of ReciNet across diverse material systems as a foundational building block.
@@ A.1 Relations Between Real and Reciprocal Lattice Vectors
In materials science, crystals exhibit translational symmetry, where atoms are periodically arranged in 3D space, forming a direct lattice in real space. Each lattice point corresponds to a repeating unit in the crystal structure, described by primitive lattice vectors $\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}$. Reciprocal space, on the other hand, represents the spatial frequencies associated with this periodic arrangement. It forms a dual lattice, mathematically related to the real lattice through reciprocal lattice vectors $\bm{b}_{1},\bm{b}_{2},\bm{b}_{3}$. $V$ is the volume of the parallelepiped spanned by the three primitive translation vectors of the original Bravais lattice. It remains consistent under periodic permutations of the indices (Economou, 2010).
@@ A.2 Diffraction Methods
Electron diffraction is widely used in transmission electron microscopy (TEM) to analyze materials. For instance, single crystals produce discrete spot patterns, polycrystals form concentric rings, and amorphous materials show diffuse rings (Egerton et al., 2005). These patterns enable crystallinity and orientation analysis, with applications in semiconductors, metals, and oxides. Figure˜3 illustrates a typical diffraction pattern. By satisfying the Bragg condition, elastically scattered electrons produce high-intensity spots, with the angular relationship between transmitted and diffracted beams revealing the crystal structure (Poeppelmeier, 2023).
@@ B.1 Dataset
JARVIS-DFT Formation Energy 55,722
Bandgap (OPT) 55,722
Total Energy 55,722
Bandgap (MBJ) 18,171
$E_{\text{hull}}$ 55,370
Materials Project Formation Energy 69,239
Band Gap 69,239
Bulk Moduli 5,450
Shear Moduli 5,450
MatBench e_form 132,752
@@
JARVIS dataset.
The JARVIS dataset, proposed by Choudhary et al. (2020), contains 55,723 materials and serves as a benchmark for crystal property prediction. Following the experimental protocols of Yan et al. (2024), we evaluate our methods on five regression tasks: formation energy, total energy, bandgap (using both OPT and MBJ methods), and energy above the convex hull (Ehull). The training, validation, and testing splits for formation energy, total energy, and OPT bandgap consist of 44,578, 5,572, and 5,572 crystals, respectively. The splits for Ehull include 44,296, 5,537, and 5,537 crystals, while the MBJ bandgap tasks consist of 14,537, 1,817, and 1,817 crystals. The MBJ functional is considered more accurate for bandgap calculations compared to OPT, and both are utilized in this study. Notably, 18,865 of the dataset’s crystal structures have been experimentally observed, adding robustness to its use in predictive modeling.
The Materials Project dataset.
The Materials Project dataset, introduced by Chen et al. (2019), is a comprehensive collection of 69,239 materials, widely utilized in crystal property prediction studies. We adopt the data splits in Yan et al. (2024) to ensure consistency and fair comparisons with prior methods. Specifically, for formation energy and bandgap prediction tasks, the dataset comprises 60,000, 5,000, and 4,239 crystals for training, validation, and testing, respectively. For bulk modulus and shear modulus predictions, the splits include 4,664, 393, and 393 crystals. Notably, one validation sample in the shear modulus task is excluded due to a negative GPa value, reflecting an underlying unstable or metastable crystal structure. Among the included crystals, 30,084 have been experimentally observed, further highlighting the dataset’s reliability for studying material properties.
MatBench dataset.
The Matbench dataset (Dunn et al., 2020) is a benchmark suite designed for evaluating machine learning models on materials science tasks, offering standardized splits and diverse property predictions. In this work, we focus on two representative regression tasks from Matbench: the large-scale formation energy prediction task (e_form) comprising 132,752 crystal structures, and the smaller-scale 2D exfoliation energy task (jdft2d) with 636 samples. These tasks assess the model’s ability to generalize across both abundant and limited data regimes.
@@ B.2 Settings of Model Embeddings
[…]
@@ B.3 Geometric GNN Information
where $h_{i}^{(l)}$ and $h_{j}^{(l)}$ are node embeddings of nodes $i$ and $j$, respectively, $d_{ij}$ is the edge attribute, and $\|$ denotes concatenation. These concatenated features are processed by a multi-layer perceptron to compute attention coefficients that regulate the influence of neighboring nodes:
@@
B.4 Hyperparameter Settings of ReciNet on Different Tasks
In this subsection, we share the detailed hyperparameter settings of ReciNet and ReciNet-MT for different tasks in JARVIS, the Materials Project, and MatBench crystal datasets. We use a different number of blocks for different tasks, and higher performance is expected if hyperparameters are tuned specifically for each task.
Table 7: Model settings of ReciNet for the JARVIS dataset.
@@ B.4 Hyperparameter Settings of ReciNet on Different Tasks
Table 8: Model settings of ReciNet for the Materials Project dataset.
@@
Table 9: Model settings of ReciNet-MT for JARVIS and the Materials Project datasets.
@@
Table 10: Model settings of ReciNet for MatBench.
@@
JARVIS.
We show the model settings of ReciNet on the JARVIS dataset in Table˜7. We train both models using the Adam (Kingma, 2014) optimizer with weight decay of 1e–5, Onecycle scheduler (Smith & Topin, 2019), and for a duration of 500 training epochs. The batch size is standardized at 64, and the models are trained using the L1 loss function. The effectiveness of the model is quantitatively measured using the mean absolute error (MAE). The number of neighbors indicates the $k$-th nearest distance we use as the radius for node $i$.
The Materials Project.
We show the model settings of ReciNet on the Materials Project dataset in Table˜8. We train both models using the Adam (Kingma, 2014) optimizer with weight decay of 1e–5, and the Onecycle scheduler (Smith & Topin, 2019). The batch size is standardized at 64, and the models are trained using L1 loss function. The effectiveness of the model is quantitatively measured using the mean absolute error (MAE). The number of training epochs for shear moduli on ReciNet is 300, while for the other three properties it is 500.
MatBench.
We show the model settings of ReciNet on the MatBench dataset in Table˜10. All models are trained for 500 epochs using the Adam (Kingma, 2014) optimizer with MAE loss. The effectiveness of the models is quantitatively measured using the mean absolute error (MAE) and root mean square error (RMSE).
B.5 Additional Ablation Studies
Table 11: Ablation study of ReciNet with varying model depth, reported in terms of test MAE on the JARVIS dataset. The numbers 3, 4, and 5 denote the number of ReciNet blocks used. The best results are shown in bold.
@@ B.5 Additional Ablation Studies
ReciNet(3) 27.8 0.127 27.2 0.242 52.9
ReciNet(4) 27.0 0.126 26.8 0.239 48.0
ReciNet(5) 27.7 0.127 27.2 0.240 43.0
@@
We further investigate the effect of model depth by varying the number of ReciNet blocks from 3 to 5. As shown in Table˜11, increasing the number of blocks from 3 to 4 consistently improves performance across all tasks. Performance on $E_{\text{hull}}$ continues to improve with five blocks, suggesting that deeper message passing may help capture the subtle energy differences relevant to thermodynamic stability. However, for other properties on the JARVIS dataset, adding a fifth block offers no additional benefit and leads to slight performance degradation. This suggests that while increased depth enhances representational capacity, it may also introduce overfitting or optimization difficulties, especially when training data is limited. To balance predictive accuracy and computational cost, we adopt three or four ReciNet blocks as the default configuration.
B.6 Additional Experimental Comparisons with Reciprocal Space-Based Methods
Table 12: Comparison of ReciNet with machine learning models incorporating reciprocal space. For EwaldMP, we use $N_{x}=N_{y}=N_{z}=1$ for simplicity. For Crystalformer, dual space means the combined use of real space and reciprocal space attention. The best results are highlighted in bold, and the second-best results are indicated with underlines.
meV/atom eV meV/atom eV meV
EwaldMP (Kosmala et al., 2023) 31.0 0.132 31.6 0.26 57
PotNet (Lin et al., 2023) 29.4 0.127 32 0.27 55
Crystalformer (Taniai et al., 2024) (dual space) 34.3 0.126 35.3 0.283 28.4
ReciNet 27.0 0.126 26.8 0.24 43
EwaldMP.
We have provided additional benchmarking results on the JARVIS dataset, where both ReciNet and EwaldMP (Kosmala et al., 2023) are evaluated. As shown in the Table˜12, ReciNet consistently outperforms EwaldMP across all five target properties, demonstrating the effectiveness of our approach beyond EwaldMP for bulk crystalline materials. EwaldMP integrated structure factor embeddings directly into GNN message passing, enabling feedback to node embeddings. However, it requires extensive hyperparameter tuning for grid sizes ($N_{x},N_{y},N_{z}$) to define the supercell, which can disrupt the symmetry of the unit cell, alter its space group, and substantially increase computational costs, particularly for systems with complex symmetries. While EwaldMP incorporates reciprocal space modeling, their approach is tailored to molecular systems and evaluated on semi-periodic surface structures from the OC20 dataset, which includes relaxed small-molecule adsorbates (e.g., CO, H2O, NH3) on metal surfaces (e.g., Pt, Cu, Ni). These systems lack full 3D periodicity and space group symmetries typical of bulk crystals. As a result, the study does not assess reciprocal space modeling in fully periodic crystalline materials such as perovskites (e.g., SrTiO3) or semiconductors (e.g., GaAs), which are typical and important materials. In contrast, our method is explicitly designed for bulk crystals and directly encodes their symmetry and periodicity.
PotNet.
PotNet (Lin et al., 2023) explicitly models long-range interactions by using the Ewald summation technique, which calculates the long-range portion of the potential in reciprocal space. However, these potentials are pre-computed and fed into the GNN as fixed, non-trainable edge features. That says, PotNet does not perform learning or message passing in reciprocal space. This limits the model’s ability to adaptively learn complex interactions from different crystalline materials. Also, PotNet incurs an additional $\mathcal{O}(n^{2})$ computational cost due to pre-calculation (which is shown in Table˜5 where the inference time of PotNet is dominated by the precomputation of potential summations). This makes it inefficient for complex material systems with large $n$, where $n$ is the number of atoms in the unit cell. In contrast, ReciNet integrates the reciprocal space transformation directly into the network’s forward pass, applying fully learnable filters to the Fourier-space representations in an end-to-end manner. The comparison with PotNet is shown in Table˜12 on the JARVIS dataset, which validates these advantages. ReciNet achieves lower MAE values for all five tasks than PotNet, especially lower MAE for formation energy (27.0 vs. 29.4), $E_{\text{hull}}$ (43 vs. 55), and total energy (26.8 vs 32). Together, these demonstrate the superiority of ReciNet compared to PotNet.
Crystalformer.
Separately, Crystalformer (Taniai et al., 2024) introduces reciprocal space via fixed analytical Fourier-Gaussian formulations within a Transformer-style architecture in their appendix. However, their reciprocal space component is non-trainable and presented only as an optional variant in Appendix J. Our proposed ReciNet instead employs fully learnable reciprocal space filters, trained end-to-end to capture long-range interactions beyond the capacity of fixed-function decays. While both models incorporate reciprocal space to enhance long-range interaction modeling, ReciNet consistently demonstrates stronger performance. As shown in Table˜12, ReciNet outperforms Crystalformer (dual space variant from Appendix J in Crystalformer) in the JARVIS dataset. Specifically, ReciNet reduces the MAE by 21.3% for formation energy (0.0270 vs. 0.0343), 24.1% for total energy (0.0268 vs. 0.0353), and 15.2% for Bandgap(MBJ) (0.24 vs. 0.283). These substantial improvements highlight the effectiveness of our learnable reciprocal space design in capturing complex, long-range dependencies in crystalline materials.